Systems Staff Meeting Notes 8/14/00

Systems Staff Meeting Notes
August 14, 2000

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Winston Harris, Bobbie Meng, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. Martha Hruska will be in charge of the Systems Department during Bill's Leave of Absence. Staff meetings will continue to be held - as will Liaison and Middle Managers. Martha will chair Middle Managers, but she has a scheduling conflict with the Liaisons timeslot. Whoever is appointed temporary chair will have to send out the meeting reminder. Systems staff will have full phone contact and access to Martha.
  2. Bill notes that there are great peripheral prices on items for home computers right now: 20 Gig HD at Best Buy - Maxtor for $120, CD Write Once - 30 Gig - $15, CD read/write 8x4x32 - about $200 (internal only). Memory and other items are available at really good prices.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Barcode readers received.
  2. Work station status: Dell station is hung up somewhere, Winston is checking on it.
  3. Bill is clearing out his old trouble tickets and he encourages everyone to do the same, as Martha likes detail.
  1. Classification Plus & Cataloguer's Desktop license expired a week early. Bill is working to fix it.
  2. North Tower - Will saw Tim on Thursday. The wiring job is on hold for some reason and Digital was anxious. Will is going to check on this.
  3. Conduit - money is now available, Facilities will purchase and install. Digital will be back - need to coordinate with Judy.
Software Projects Status
  1. Employee-Patron Update blew up. A special request to Hilton was to take the 9 & 10 month faculty off the list. Tigert said the list should be cleaned up this week. Exchange/Access Services is still doing some A&P manually requested faculty status.
  2. Dean's Menu - will be handled by Suzy & Debra in Bill's absence. Dean's Menu relates to other things - CMS may go away, which will affect how we do our batch submissions. It is scheduled to go away on the first of March - we'll try to move utilities to Unix. TSO is too expensive. We received a request to cancel Ed Teague's VM account from NERDC.
  3. Ying had a meeting with Erich & Stephanie - they have many new requests. Ying & Winston may reschedule their meetings with Erich for every other Tuesday instead of every Tuesday.
DP Coordinator
  1. Items sent to West via mail - leave slips, packing slips, etc. - seem to not be arriving at their intended destinations. Bill will investigate.
  2. All students except for Jorga are out this week.
  1. Debra out today & tomorrow.
  2. Bill's last day will be August 22.
  3. Labor Day Holiday - Mon. Sept. 4.

Back to the Systems Homepage

Last updated August 22, 2000
by Debra Harris